# Exploring Lattice Thermal Conductivity Models via Interpretable Deep Learning to Accelerate Materials Discovery

### v7

@@ II.2 Deepening physical interpretability through sensitivity analysis
$\log{\left(\kappa\right)_{{\rm{KAN}}}}=-1.05\sin{f_{1}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)+0.61\cosh{f_{2}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)-1.44+1.4{{\rm{e}}^{-19.72f_{3}^{2}\left({F,{K_{s}},\alpha,{C_{v}}}\right)}}-0.48{{\rm{e}}^{-63.21f_{4}^{2}\left({{C_{v}},{K_{s}},F}\right)}}$ (5)
@@
${f_{1}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)$ $=-5.66{\left({0.77-F}\right)^{2}}-0.7\tan\left({1.56{K_{s}}-4.4}\right)+2.89+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{v}}-0.11}\right)}^{2}}}},$
${f_{2}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)$ $=-0.76\sin\left({4.23F+4.79}\right)+1.03\tan\left({1.55\alpha-0.99}\right)-0.86\cosh\left({5.35{C_{v}}-1.75}\right)$
$\quad+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{s}}-0.63}\right)+2.31,$
${f_{3}}\left({F,{K_{s}},\alpha,{C_{v}}}\right)$ $={\left({0.38-{K_{s}}}\right)^{3}}-0.09\sin\left({6.2{C_{v}}-6.16}\right)-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)+0.47-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}},$
${f_{4}}\left({{C_{v}},{K_{s}},F}\right)$ $=-{\left({0.37-{C_{v}}}\right)^{2}}-0.13\tan\left({1.46{K_{s}}-0.63}\right)-0.3+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}.$
@@
[…]
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Figure 4(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between LTC and the features, while also demonstrating exceptional robustness.
@@ II.3 Screening potential thermal insulators/conductors
Unlike simple descriptors that can be directly generated by programs, such as those based on composition (e.g., Magpie [67]) or structure (e.g., Coulomb matrix [68]), the physical features involved in the aforementioned reduced dataset are relatively complex and not easily constructed physical quantities. Inspired by previous studies [69, 70], our approach is to use a two-stage prediction method. We predict features by simple descriptors and models, and then use the predicted features to perform high-throughput predictions of the LTC for unknown materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties [35]. In our application, CGCNNs delivered fairly accurate results for the prediction of all features except for $C_{\text{v}}$, where the prediction accuracy was notably lower (see Supplementary Information for details). Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using features that can be reliably predicted qualitatively, and then calculating $C_{\text{v}}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate LTC.
Figure 5: Correlation between the 4 features and LTC. (a)-(d) The two-dimensional mapping of different features with $\log(\kappa)$, where the color bar reflects the scatter density, and the feature values have been min-max normalized. (e) The mapping between $\log(\kappa)$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa)$ of each sample. The dashed line represents the mean $\log(\kappa)$ value of 1 for the samples.
Figure 6: Using CGCNN to predict key features and qualitatively assess LTC. (a), (b), and (c) display scatter plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets. (d), (e), and (f) illustrate the distribution curves of $F$, $K_{\text{s}}$, and scores for Materials Project samples, ordered by their comprehensive scores, with the red and blue scatter points of varying shapes representing the top 5 and bottom 5 samples, respectively.
Based on the information from Figure 3(a) and Figure 5(a)-(d), it can be concluded that $F$, $K_{\text{s}}$, and $\kappa$ exhibit a strong positive linear correlation. Materials with higher F and $K_{\text{s}}$ values generally exhibit higher LTC, a relationship that is further confirmed in Figure 5(e). Therefore, the preliminary qualitative screening of thermal insulators and conductors based on F and $K_{\text{s}}$ is highly beneficial for accelerating the exploration of new materials. The predictions of F and $K_{\text{s}}$ using CGCNN are shown in Figure 6(a) and 6(b), with $R^{2}$ values reaching 0.94 and 0.98, respectively. We extracted a total of 2,246 samples from the Materials Project [71] with the following selection criteria:
@@
Additionally, we developed a comprehensive score to qualitatively evaluate a sample’s potential to become a thermal insulator or conductor. This score incorporates the linear coefficients between features and targets, the Kucherenko indices of features with respect to the target, and the features themselves:
${\rm{Score}}=\frac{{{K_{F}}}}{{{K_{F}}+{K_{{K_{s}}}}}}\frac{{{R_{F}}}}{{{R_{F}}+{R_{{K_{s}}}}}}F+\frac{{{K_{{K_{s}}}}}}{{{K_{F}}+{K_{{K_{s}}}}}}\frac{{{R_{{K_{s}}}}}}{{{R_{F}}+{R_{{K_{s}}}}}}{K_{s}}$ (7)
where $K$ represents the respective Kucherenko index, and $R$ denotes the Pearson correlation coefficient. Based on the scores of the materials, we retained the top 5 and bottom 5 samples to verify their potential as thermal conductors or insulators. Existing literature [72, 73, 74, 75, 76, 77] on these samples primarily focuses on 2D materials or electronic thermal conductivity. In contrast, our research targets the LTC of 3D materials, aiming to fill this gap in the field.
@@
Figure 7: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the thermal conductivity to converge. The deep red dashed line represents the thermal conductivity predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The LTC calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) shows a histogram of the thermal conductivity distribution in the original dataset.
Based on the conclusions provided by the MP database, most of the 10 selected structures are predicted to be unstable; consequently, only a few materials, such as CN2, Be2C, CrB4, CuBO2, and BeSiO3, possess valid ${C_{\rm{v}}}$ values obtained through DFT calculations (see Table 3). We utilized GPUMD to calculate both the surface and bulk thermal conductivities of viable candidate materials. For isotropic crystals, we computed results in only one direction. For anisotropic crystals, we calculated the LTC in each direction and averaged these values to represent the overall performance. Notably, CN2 exhibits thermal conductivity exclusively in the $z$-direction. As such, averaging the LTC across multiple directions for this material is not meaningful, and the z-direction thermal conductivity is used to represent its overall performance. Figure 7(a), (b), and (c) illustrate the LTC convergence for each material. We calculated both the surface and bulk thermal conductivities of each material over relaxation times ranging from 0 to 10 ns, providing the sum of these values for comparison with the predictions from the KAN model. In our work, the pretraining model faced several challenges:
  • Feature selection: Initially, the original dataset provided 10-dimensional features for training all models. However, to reduce model complexity, enhance interpretability, and account for the difficulty of obtaining certain features, we applied sensitivity analysis to narrow the features down to just four dimensions. While both KAN and the black-box models demonstrated remarkable robustness, the uncertainty introduced by feature selection remains a consideration.
  • Two-stage prediction: As shown in Figure 6, three features for the candidate materials were derived from CGCNN predictions, with only $C_{\text{v}}$ obtained from DFT calculations. Although CGCNN and DFT provide satisfactory accuracy for individual properties, the two-stage prediction process inevitably introduces cumulative errors in predicting the LTC of candidate materials.
  • Data imbalance [78]: As depicted in Figure 7(d), only 0.47% of the samples in the dataset used for model pretraining have an LTC $>$ 100 W/m/K. Consequently, the model tends to favor the more abundant low- and medium-conductivity materials, potentially limiting its extrapolation capability in unrepresented regions.
According to our MD calculations, the candidate materials exhibit LTC values within distinct ranges (3̃0, 8̃0, 1̃50 W/m/K). Despite this, their ML-predicted LTC consistently align with expectations. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability. Although the qualitative scores for CrB4 and Be2C are quite close, there is a significant difference in their actual thermal conductivities. This discrepancy is attributed to a notable difference in their heat capacities (see Table 3), which is not accounted for in Eq. 6 due to its negligible negative correlation with LTC. However, overall, the qualitative assessment method in this study helped us identify other new materials with excellent thermal conductivity, and it proved valuable in improving the efficiency of material screening.
Table 3: The predicted results of LTC based on KAN.
MPID Formula $F_{pred}$ $K_{s-pred}$ $\gamma_{pred}$ $C_{v-DFT}$ $\kappa_{KAN}$
mp-1009818 CN2 152.8788 333.6628 $1.71^{-5}$ 0.2343 155.2907
mp-27710 CrB4 113.7286 259.3992 $2.73^{-5}$ 12.7181 29.3775
mp-1569 Be2C 138.8222 191.1319 $2.41^{-5}$ 4.6692 80.5562
mp-1096940 CuBO2 78.15269 267.6422 $3.04^{-5}$ 8.4535 31.3254
mp-1183445 BeSiO3 85.25379 224.8873 $2.56^{-5}$ 9.5514 41.3833
mp-1095948 Cs2KNa -139.976 18.569 / / /
mp-1097263 Cs2RbNa -144.966 16.8264 / / /
mp-10378 Cs3Sb -143.5 11.2367 / / /
mp-1097633 Cs2KRb -145.791 14.89677 / / /
mp-635413 Cs3Bi -168.671 13.74844 / / /
@@ III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the LTC of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based LTC modeling has achieved remarkable accuracy, but inprovments in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of LTC modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
In this study, we employed KAN to model LTC, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features, highlighting its superior interpretability in this context. Additionally, the analytical expressions derived from KAN serve as a valuable supplement to traditional empirical models. We also developed a two-stage high-throughput prediction framework, integrating CGCNN to accurately predict the primary physical features required for LTC. For $C_{\text{v}}$, a property that CGCNN cannot predict accurately, DFT calculations were used. Building on this foundation, the predictions from the KAN pretraining model aligned closely with MD validation results, enabling us to successfully identify three new materials, one of which was confirmed as an excellent thermal conductor. Compared to using MD to directly compute LTC for unknown materials, our high-throughput prediction framework requires only DFT calculations of $C_{\text{v}}$ to achieve MD-level predictive accuracy. This work not only accelerates the prediction of material thermal conductivity but also enhances our understanding of the underlying physical mechanisms. By balancing accuracy and interpretability, KAN demonstrates its potential as a powerful tool for advancing materials informatics.

### v8

@@ II.2 Deepening physical interpretability through sensitivity analysis
$\log{\left(\kappa\right)_{{\rm{KAN}}}}=$ $-1.05\sin d_{1}+0.61\cosh d_{2}-1.44$
$+1.4{{\rm{e}}^{-19.72d_{3}^{2}}}-0.48{{\rm{e}}^{-63.21d_{4}^{2}}}$ (5)
@@
$d_{1}$ $=-5.66{\left({0.77-F}\right)^{2}}-0.7\tan\left({1.56{K_{\rm s}}-4.4}\right)+2.89$
$\quad+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{\rm v}}-0.11}\right)}^{2}}}}$
$d_{2}$ $=-0.76\sin\left({4.23F+4.79}\right)+1.03\tan\left({1.55\alpha-0.99}\right)$
$\quad-0.86\cosh\left({5.35{C_{\rm v}}-1.75}\right)+2.31$
$\quad+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{\rm s}}-0.63}\right)$
$d_{3}$ $={\left({0.38-{K_{\rm s}}}\right)^{3}}-0.09\sin\left({6.2{C_{\rm v}}-6.16}\right)$
$\quad-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)+0.47$
$\quad-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}}$
$d_{4}$ $=-{\left({0.37-{C_{\rm v}}}\right)^{2}}-0.13\tan\left({1.46{K_{\rm s}}-0.63}\right)-0.3$
$\quad+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}$
@@
$\log{\left(\kappa\right)_{{\rm{SISSO}}}}=$ $1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}$
$-45.57\sqrt[3]{{{\alpha^{2}}{C_{\rm{v}}}{K_{\rm{s}}}}}$ (6)
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Figure 4(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between $\kappa_{\rm L}$ and the features, while also demonstrating exceptional robustness.
@@ II.3 Screening potential thermal insulators/conductors
Unlike simple descriptors that can be directly generated by programs, such as those based on composition (e.g., Magpie [67]) or structure (e.g., Coulomb matrix [68]), the physical features involved in the aforementioned reduced dataset are relatively complex and not easily constructed physical quantities. Inspired by previous studies [69, 70], our approach is to use a two-stage prediction method. We predict primary physical features by CGCNNs, and then use the predicted primary features to perform high-throughput predictions of the $\kappa_{\rm L}$ for unlabeled materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties [35]. In our application, CGCNNs delivered fairly accurate results for the prediction of all primary features except for $C_{\text{v}}$, where the prediction accuracy was notably lower (see Supplementary Information for details).Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using primary features that can be reliably predicted qualitatively, and then calculating $C_{\text{v}}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate $\kappa_{\rm L}$.
Figure 5: Predicting key features and qualitatively assessing $\kappa_{\rm L}$ via CGCNN. (a), (b), and (c) display scatter plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets, where $R^{2}_{F}=0.940$, $R^{2}_{K_{\rm s}}=0.980$ and $R^{2}_{\alpha}=0.910$.
Based on the information from Figure 3(a) and Figure 8(a)-(d), it can be concluded that $\kappa_{\rm L}$ is strongly and positively linearly correlated with both $F$ and $K_{\rm s}$. Materials with higher F and $K_{\text{s}}$ values generally exhibit higher $\kappa_{\rm L}$, a relationship that is further confirmed in Figure 8(e). Therefore, the preliminary qualitative screening of thermal insulators and conductors based on $F$ and $K_{\text{s}}$ is highly beneficial for accelerating the exploration of new materials. The predictions of $F$, $K_{\text{s}}$ and $\alpha$ using CGCNN are shown in Figure 5(a)-(c), with $R^{2}$ values reaching 0.94, 0.98 and 0.91, respectively. We extracted a total of 2,246 samples from the Materials Project (MP) [71] with the following selection criteria:
@@
Although CGCNN is capable of accurately evaluating $F$, $K_{\rm s}$, and $\alpha$ for unlabeled materials, to the best of our knowledge, there is currently no reliable machine learning model that exhibits outstanding performance in predicting $C_{\rm v}$. When performing a quantitative estimation of $\kappa_{\rm L}$ for a large number of new materials based on Eq. (5), the $C_{\rm v}$ parameter still needs to be obtained from DFT calculations. However, this approach imposes a considerable computational cost when dealing with a large number of samples. To address this, we carry out unsupervised learning to rapidly screen the most promising thermal conductors and insulators from the MP database, enabling precise identification of candidate materials.
Figure 6: $k$-means clustering and t-SNE 2D visualization. (a) Labeled training samples from AFLOW are clustered into seven groups, each represented by a different color. (b) The same scatter plot as in (a), but the samples are colored according to their $\log(\kappa_{\rm L})$, with black pentagrams representing unlabeled candidates from the MP database. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the t-SNE space.
For this purpose, we first attempt to use $k$-means clustering [72] on samples from the AFLOW database (e.g., the labeled dataset used for model training in Section II.1) to evaluate its effectiveness. Since reliable quantitative estimation is currently only feasible for the three primary features $F$, $K_{\rm s}$, and $\alpha$, we retain only these three features for each sample during clustering. In $k$-means clustering, the optimal number of clusters is typically determined using the elbow method [73] or the silhouette score [74]. However, in our experiments, these two methods yielded inconsistent results, making it unclear which one is more convincing (see Supplementary Information). Fortunately, through manual inspection, we found that when $K=4$, the clustering results already exhibit a clear and meaningful trend.
In Figure 6(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Figure 6(a), as shown in Figure 6(b), we observe that as t-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Figure 6(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 6(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Therefore, we focus on these 10 candidates for further predictions and validation. Existing literature [75, 76, 77, 78, 79, 80] on these candidates primarily focuses on 2D materials or electronic thermal conductivity. Our research targets the $\kappa_{\rm L}$ of 3D materials, aiming to fill this gap in the field.
@@
Figure 7: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the thermal conductivity to converge. The deep red dashed line represents the thermal conductivity predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The $\kappa_{\rm L}$ calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) shows a histogram of the thermal conductivity distribution in the primary dataset.
Figure 8: Correlation between the 4 key primary features and $\kappa_{\rm L}$. (a)-(d) Two-dimensional mappings of different features with $\log(\kappa_{\rm L})$, where the feature values have been min-max normalized. The closed solid lines represent contour lines of the sample density distribution, while the dashed lines indicate fifth-order polynomial fitting curves. (e) The mapping between $\log(\kappa_{\rm L})$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa_{\rm L})$ of each sample. The dashed line represents the mean $\log(\kappa_{\rm L})$ value of 1 for the samples.
According to our MD calculations, the actual $\kappa_{\rm L}$ values of the candidate materials fall within distinct ranges ($\sim$30, $\sim$80, $\sim$150 W/m/K). As shown in Figure 7(d), samples with $\kappa_{\rm L}>100$ W/m/K account for only 0.47% of the entire training dataset. In such cases of data imbalance, model predictions for the target region with scarce samples typically exhibit significant uncertainty [81, 82]. Nevertheless, the $\kappa_{\rm L}$ values predicted by KAN remain in excellent agreement with the MD validation results. This demonstrates that the pretrained KAN model achieves not only reliable accuracy but also satisfactory extrapolation capability. The methods of $k$-means clustering and t-SNE visualization are also sufficient for qualitatively assessing the $\kappa_{\rm L}$ of materials, enabling high-throughput screening in combination with the pretrained model.
Table 3: The predicted results of $\kappa_{\rm L}$ based on KAN
MPID Formula $F_{\rm pred}$ $K_{\rm s-pred}$ $\alpha_{\rm pred}$ $C_{\rm v-DFT}$ $\kappa_{\rm L-KAN}$
mp-1009818 CN2 152.8788 333.6628 $1.71\times 10^{-5}$ 0.2343 155.2907
mp-27710 CrB4 113.7286 259.3992 $2.73\times 10^{-5}$ 12.7181 29.3775
mp-1569 Be2C 138.8222 191.1319 $2.41\times 10^{-5}$ 4.6692 80.5562
mp-1096940 CuBO2 78.15269 267.6422 $3.04\times 10^{-5}$ 8.4535 31.3254
mp-1183445 BeSiO3 85.25379 224.8873 $2.56\times 10^{-5}$ 9.5514 41.3833
mp-1095948 Cs2KNa -139.976 18.569 $4.84\times 10^{-4}$ / /
mp-1097263 Cs2RbNa -144.966 16.8264 $4.90\times 10^{-4}$ / /
mp-10378 Cs3Sb -143.5 11.2367 $2.42\times 10^{-4}$ / /
mp-1097633 Cs2KRb -145.791 14.89677 $5.10\times 10^{-4}$ / /
mp-635413 Cs3Bi -168.671 13.74844 $2.50\times 10^{-4}$ / /
II.4 Tracing high/low $\kappa_{\rm L}$ from a theoretical perspective
@@ III DISCUSSION
In the past, researchers often relied on empirical or semi-empirical models to calculate the $\kappa_{\rm L}$ of materials. However, as the range of explored materials continues to expand, the inaccuracies inherent in empirical models have become increasingly problematic. Machine learning-based $\kappa_{\rm L}$ modeling has achieved remarkable accuracy, but inprovments in accuracy alone offers limited contributions to the advancement of materials science. Black-box models, such as neural networks and ensemble learning, excel in accuracy but their complex structures hinder the understanding of feature-target relationships. In contrast, white-box models like symbolic regression provide transparent structures, making their internal mechanisms interpretable, but this often comes at the cost of accuracy. Traditionally, the interpretability and accuracy of ML or DL models have been seen as mutually exclusive, akin to “having your cake and eating it too.” However, in the context of $\kappa_{\rm L}$ modeling, interpretable DL models like KAN have demonstrated significant success in balancing both accuracy and interpretability.
In this study, we employed KAN to model $\kappa_{\rm L}$, demonstrating that its performance in terms of both accuracy and robustness is fully comparable to that of black-box models. For the interpretability analysis of white-box models, we considered feature interaction effects. Sensitivity analysis results based on the KAN and SISSO models revealed that KAN can accurately extract key features, highlighting its superior interpretability in this context. Additionally, the analytical expressions derived from KAN serve as a valuable supplement to traditional empirical models. We also developed a two-stage high-throughput prediction framework, integrating CGCNN to accurately predict the primary physical features required for $\kappa_{\rm L}$. For $C_{\text{v}}$, a property that CGCNN cannot predict accurately, DFT calculations were used. Building on this foundation, the predictions from the KAN pretraining model aligned closely with MD validation results, enabling us to successfully identify three new materials, one of which was confirmed as an excellent thermal conductor. Compared to using MD to directly compute $\kappa_{\rm L}$ for unknown materials, our high-throughput prediction framework requires only DFT calculations of $C_{\text{v}}$ to achieve MD-level predictive accuracy. This work not only accelerates the prediction of material thermal conductivity but also enhances our understanding of the underlying physical mechanisms. By balancing accuracy and interpretability, KAN demonstrates its potential as a powerful tool for advancing materials informatics.